Suppose there is a McDonald's next to your school and you want to estimate the proportion, p of all people at your school who go to McDonald's on a typical school day.

(a) If you had no preliminary estimate for p. how many people should you include in a random sample to be 85% sure that the estimate will be within a distance of 0.05 from p? Make sure to show the formula with the numbers plugged in.

(b) Complete the directions in part A knowing that about 1 in 20 students go to McDonald's every day. Make sure to show the formula with the numbers plugged in.

(c) If you want a smaller margin of error, what will you need to do to the sample size?
Answer #1

Solution:-

a) The number of people we should include in a random sample to be 85% sure that the estimate will be within a distance of 0.05 from p is 208.

M.E = 0.05, C.L = 85%

p = 0.50 (No preliminary estimate)

p(1 - p)

n =left ( rac{z_{alpha /2}}{M.E} ight )^{2} imes p(1-p)

1.44 0.05 × 0.50(1-.0.50)

n = 207.36

n = 208

b) The number of people we should include in a random sample to be 85% sure that the estimate will be within a distance of 0.05 from p is 40.

M.E = 0.05, C.L = 85%

p = 1/20 = 0.05 (Given)

p(1 - p)

n =left ( rac{z_{alpha /2}}{M.E} ight )^{2} imes p(1-p)

1.44 005 × 0.05(1-0.05)

n = 39.39

n = 40

c) For smaller margin of error we need to increase the sample size.

p(1 - p)